# Beyond Random: Automatic Inner-loop Optimization in Dataset Distillation

### v1

@@ 5.4 Ablation Study
Component Combination Evaluation. In this section, we evaluate the individual and combined contributions of three performance components: dynamic truncation position (DTP), adaptive window size (AWS), and patch-wise semantic preservation (PSP). Ablation studies are conducted under the IPC=10 setting across CIFAR-10/100 krizhevsky2009learning , Tiny-ImageNet le2015tiny and ImageNet-1K DBLP:journals/ijcv/RussakovskyDSKS15 . Tab. 3 shows that the integrated application of DTP and AWS yields a combined optimization effect, resulting in a 2.8% accuracy improvement, which surpasses the sum of the individual contributions. This arises from the complementary information processing mechanisms of DTP and AWS, which enhance the model’s ability to capture key features of the original dataset. Regarding the PSP component, Tab. 3 reveals that its accuracy improvement is marginal on CIFAR-10/100 yet it is substantial on two datasets with higher resolution. When combined with DTP and AWS, it leads to remarkable gains of 8.3% and 17.6%. This is attributed to PSP’s ability to segment high-resolution images into smaller patches, thereby fully leveraging AT-BPTT’s strengths across various resolutions.
@@
Stage Transition Threshold Analysis. To achieve precise control over stage transitions during model training, we study four critical hyperparameters: $M$ (the accuracy variation threshold in the early stage), $X$ (the counter threshold for early stage), $N$ (the accuracy variation threshold in the middle stage), and $Y$ (the counter threshold for middle stage). Through experiments with the CIFAR-10 krizhevsky2009learning under IPC=10, we identify that suboptimal parameter configurations induce two failure modes: the plateaus in the initial training phase or too early transition to the final stage, both of which disrupt the staged paradigm and cause substantial performance degradation. Fig. 5 demonstrates optimal parameterization when $X$ corresponds to 5% of total training epochs and $Y$ to 4%, with MM=1.5 and NN=1.0. This configuration achieves balanced stage duration distribution, stable convergence patterns, and better AT-BPTT performance.
@@
Window Size Analysis. The hyperparameter $d$ in Section 4.2 governs the adjustment range of the truncation window, demonstrating a trade-off between model performance and computational efficiency. We conduct an ablation study on CIFAR-10 krizhevsky2009learning with IPC=10 and quantitatively assess the dual effects of varying $d$ values. As Tab.3 shows, progressive increases in $d$ simultaneously enhance test accuracy and computational demands. Notably, compared to the baseline ($d=0$), $d=10$ achieves a 1.88% accuracy improvement with modest resource increments, with only 0.6 hours additional training time and 0.8 GB GPU memory usage. However, when $d$ >> 10, the accuracy gains diminish while the computational cost rises substantially. This mainly stems from the need for longer timesteps and the corresponding exponential growth in the dimensionality of the Hessian matrix as $d$ increases. Based on a trade-off analysis between performance gains and computational costs, we set the value of $d$ to 10 and report the results obtained with this setting. More ablation studies are provided in Appendix C.
@@ Appendix B More Implementation details
To validate the effectiveness of our method while ensuring experimental rigor, we employ a controlled variable approach in designing the experimental protocol. The critical experimental parameters and configurations are rigorously aligned with those of RaT-BPTT [9], specifically encompassing essential aspects including data preprocessing procedures, model convergence acceleration mechanisms, gradient explosion mitigation strategies, and label processing methodologies. Each reported result is the average of 5 experiments..
@@ C.3 Generalization Experiments on Wide Networks
C.4 Cross-architecture Performance Evaluation
Under the IPC=10 and IPC=50 configurations on CIFAR-10 [16], we conduct cross-architecture evaluations of the distilled dataset obtained through distillation. The comprehensive evaluation spans four distinct architectures: ConvNet [11], VGG11 [36], AlexNet [17], and ResNet18 [13], accompanied by systematic comparisons with prevailing methods. As demonstrated in Tab.7, AT-BPTT consistently outperforms all baseline methods across four tested architectures, establishing new SOTA benchmarks. These findings not only validate AT-BPTT’s remarkable cross-architecture transfer capability but also provide compelling evidence for its superior generalization properties.
Table 7: Comparison with baselines across different architectures. Under the settings of IPC=10 and IPC=50 on CIFAR-10 [16], the distilled datasets are evaluated for performance on ConvNet, VGG11, AlexNet, and ResNet18. The best results are highlighted in bold.
Architecture ConvNet VGG11 AlexNet ResNet18
Img/class(IPC) 10 50 10 50 10 50 10 50
KIP [31] 47.6 42.1 24.4 36.8
DSA [55] 52.1 43.2 35.9 42.8
MTT [2] 64.3 71.6 50.3 55.4 34.2 48.2 46.4 61.9
FTD [8] 66.1 73.8 47.0 58.4 35.9 53.8 53.2 65.7
ATT [24] 67.7 74.5 54.2 61.7 43.6 60.0 54.5 66.3
NCFM [47] 71.8 77.4 68.0 75.5 67.9 75.5 67.7 73.8
Ours 72.3 78.9 68.9 77.1 68.5 76.3 69.3 75.6
C.5 High IPC Evaluation
We conduct performance comparisons between AT-BPTT and other dataset distillation methods under high IPC settings across various datasets, including CIFAR-10, CIFAR-100, and Tiny-ImageNet. Consistent with the experimental setups of other methods for high IPC, we select ResNet-18 as the evaluation network. To maintain a consistent compression ratio across datasets of different scales, we adjust the IPC settings accordingly and choose DSA [57], DM [56], MTT [2], SRe2L [51], DATM [12], SelMatch [19], CDA [50], CUDD [28], and CCFS [4] as baselines. The results shown in the Tab. 8 indicate that a significant portion of distillation methods gradually lose effectiveness as IPC increases, with performance occasionally inferior to random selection. By employing distinct distillation strategies for different stages and directly capturing model training dynamics in the inner loop, AT-BPTT effectively captures and retains deeper feature information. Consequently, even under high IPC settings, AT-BPTT outperforms other distillation methods, achieving state-of-the-art distillation performance. Notably, for Tiny-ImageNet at IPC=100, AT-BPTT attains a test accuracy of 60.7%, closely approaching the 60.9% test accuracy of the full dataset, achieving lossless distillation.
Table 8: Comparison with various baseline methods under high IPC settings using CIFAR-10/100 and Tiny-ImageNet, evaluated with the ResNet-18 network architecture, with bolded values indicating the highest test accuracy. Entries marked with “-” missing due to scalability Issues.
Dataset CIFAR-10 CIFAR-100 Tiny-ImageNet
Img/class(IPC) 250 500 1000 1500 25 50 100 150 50 100
Ratio 5% 10% 20% 30% 5% 10% 20% 30% 10% 20%
Random [57] 73.4 79.3 85.6 88.3 35.8 40.7 53.2 60.3 30.1 40.1
DSA [57] 74.7 78.7 84.8 - 38.4 43.6 - - 27.8 -
DM [56] 75.3 79.1 85.6 - 37.5 42.6 - - 31.0 -
MTT [2] 80.7 82.2 86.1 88.6 49.9 51.3 58.7 63.1 40.3 44.2
SRe2L [57] 77.5 85.1 86.8 87.8 49.7 51.4 58.8 61.9 41.1 49.7
DATM [12] - 84.8 87.6 - - 51.0 61.5 - 42.2 -
SelMatch [19] 82.8 85.9 90.4 91.3 50.9 54.5 62.4 67.4 44.7 50.4
CDA [50] 78.0 84.4 86.4 87.5 50.6 59.7 61.1 63.4 45.6 52.4
CUDD [28] - - - - 63.5 65.7 - - 55.6 56.8
CCFS [4] 87.9 92.5 93.2 93.8 65.3 71.5 73.0 74.8 55.8 60.2
Ours 89.4 93.3 94.9 95.3 66.1 72.9 75.8 76.5 57.3 60.7
Full dataset 95.5 78.8 60.9
C.6 Comparison with Diffusion Model-Based Methods
@@
C.7 Comparison with Baseline on Other Datasets
@@
C.8 Experiments on Language Tasks
@@ Appendix F Visualizations of Distilled Dataset
In this section, we present visualizations of the distilled datasets obtained from various datasets. The IPC of the presented datasets is uniformly set to 10, with Fig. 7 and Fig. 8 illustrating the visualization results on CIFAR-10 and CIFAR-100, respectively. For CIFAR-100, we display a subset of the content to ensure clarity in visualization.
@@
Figure 8: Visualization results of our proposed method for CIFAR-100 [16] with IPC=10.

### v2

@@ 5.4 Ablation Study
Component Combination Evaluation. In this section, we evaluate the individual and combined contributions of three performance components: dynamic truncation position (DTP), adaptive window size (AWS), and patch-wise semantic preservation (PSP). Ablation studies are conducted under the IPC=10 setting across CIFAR-10/100 Krizhevsky et al. (2009), Tiny-ImageNet Le and Yang (2015) and ImageNet-1K Russakovsky et al. (2015). Tab. 3 shows that the integrated application of DTP and AWS yields a combined optimization effect, resulting in a 2.8% accuracy improvement, which surpasses the sum of the individual contributions. This arises from the complementary information processing mechanisms of DTP and AWS, which enhance the model’s ability to capture key features of the original dataset. Regarding the PSP component, Tab. 3 reveals that its accuracy improvement is marginal on CIFAR-10/100 yet it is substantial on two datasets with higher resolution. When combined with DTP and AWS, it leads to remarkable gains of 8.3% and 17.6%. This is attributed to PSP’s ability to segment high-resolution images into smaller patches, thereby fully leveraging AT-BPTT’s strengths across various resolutions.
@@
Stage Transition Threshold Analysis. To achieve precise control over stage transitions during model training, we study four critical hyperparameters: $M$ (the accuracy variation threshold in the early stage), $X$ (the counter threshold for early stage), $N$ (the accuracy variation threshold in the middle stage), and $Y$ (the counter threshold for middle stage). Through experiments with the CIFAR-10 Krizhevsky et al. (2009) under IPC=10, we identify that suboptimal parameter configurations induce two failure modes: the plateaus in the initial training phase or too early transition to the final stage, both of which disrupt the staged paradigm and cause substantial performance degradation. Fig. 5 demonstrates optimal parameterization when $X$ corresponds to 5% of total training epochs and $Y$ to 4%, with MM=1.5 and NN=1.0. This configuration achieves balanced stage duration distribution, stable convergence patterns, and better AT-BPTT performance.
@@
Window Size Analysis. The hyperparameter $d$ in Section 4.2 governs the adjustment range of the truncation window, demonstrating a trade-off between model performance and computational efficiency. We conduct an ablation study on CIFAR-10 Krizhevsky et al. (2009) with IPC=10 and quantitatively assess the dual effects of varying $d$ values. As Tab.3 shows, progressive increases in $d$ simultaneously enhance test accuracy and computational demands. Notably, compared to the baseline ($d=0$), $d=10$ achieves a 1.88% accuracy improvement with modest resource increments, with only 0.6 hours additional training time and 0.8 GB GPU memory usage. However, when $d$ >> 10, the accuracy gains diminish while the computational cost rises substantially. This mainly stems from the need for longer timesteps and the corresponding exponential growth in the dimensionality of the Hessian matrix as $d$ increases. Based on a trade-off analysis between performance gains and computational costs, we set the value of $d$ to 10 and report the results obtained with this setting. More ablation studies are provided in Appendix C.
@@ Appendix B More Implementation details
To validate the effectiveness of our method while ensuring experimental rigor, we employ a controlled variable approach in designing the experimental protocol. The critical experimental parameters and configurations are rigorously aligned with those of RaT-BPTT [9], specifically encompassing essential aspects including data preprocessing procedures, model convergence acceleration mechanisms, gradient explosion mitigation strategies, and label processing methodologies. Each reported result is the average of 5 experiments.
@@ C.3 Generalization Experiments on Wide Networks
C.4 Comparison with Diffusion Model-Based Methods
@@
C.5 Comparison with Baseline on Other Datasets
@@
C.6 Experiments on Language Tasks
@@ Appendix F Visualizations of Distilled Dataset
In this section, we present visualizations of the distilled datasets obtained from various datasets. The IPC of the presented datasets is uniformly set to 10, with Fig. 7 illustrating the visualization results on CIFAR-10, respectively.